OpenStudy (anonymous):

how do I find the point-slope equation for (1,-9)

OpenStudy (anonymous):

y-y1=m(x-x1) is point slope equation

OpenStudy (anonymous):

What does it mean when it asks for the slope intercept form

OpenStudy (anonymous):

after you plug all that in, it should be y+9=(whatever the slope is)(x-1) then you solve for y